# Part-aware Shape Generation with Latent 3D Diffusion of Neural Voxel Fields

### v2

@@ Abstract
This paper presents a novel latent 3D diffusion model for the generation of neural voxel fields, aiming to achieve accurate part-aware structures. Compared to existing methods, there are two key designs to ensure high-quality and accurate part-aware generation. On one hand, we introduce a latent 3D diffusion process for neural voxel fields, enabling generation at significantly higher resolutions that can accurately capture rich textural and geometric details. On the other hand, a part-aware shape decoder is introduced to integrate the part codes into the neural voxel fields, guiding the accurate part decomposition and producing high-quality rendering results. Through extensive experimentation and comparisons with state-of-the-art methods, we evaluate our approach across four different classes of data. The results demonstrate the superior generative capabilities of our proposed method in part-aware shape generation, outperforming existing state-of-the-art methods.
Keywords:
Shape Generation Diffusion Model Part-aware Generation
1 Introduction
@@
Recent generative models have emerged for generating 3D shapes with different representations [10, 44, 37, 14], among which neural field representation [4, 3, 12, 27, 33, 46] receives increasing attention due to many advantages such as integrated texture and shape generation. Despite the fast and notable progress made along this line of research, what has been largely missed is part/structure awareness. Most existing generative models based on neural fields are part-oblivious. They tend to generate 3D shapes in a holistic manner, without comprehending its compositional parts explicitly.
@@
Generating 3D shapes with part information facilitates several downstream processing/tasks, such as editing, mix-and-match modeling, and segmentation learning. SPAGHETTI [15] and Part-NeRF [42] are two excellent part-aware generative models for 3D shape synthesis with neural field representation. Despite their strengths, there is still potential for enhancing the generative ability. SPAGHETTI relies on 3D supervision and mainly consider the geometry, ignoring the textured shape generation. On the other hand, Part-NeRF requires only 2D supervision but adopts a complex architecture with multiple neural radiance fields, leading to increased complexity in learning and subpar rendering outcomes. Moreover, both models perform object segmentation in an unsupervised manner, resulting in structurally insignificant information.
@@
We propose a latent 3D diffusion model to generate neural voxel fields with accurate structural details in this paper. As shown in Fig 1, our network is composed of two parts, a latent 3D diffusion model of neural voxel fields and a part-aware shape decoder. Specifically, the latent 3D diffusion model utilizes the advantage of 3D convolution and employs the 3D denoising process in a latent space with much lower resolution. On the other hand, the part-aware shape decoder incorporates the part code with the neural voxel field, guiding the precise generation of part-aware shapes and synthesizing high-quality rendering results.
@@
Benefiting from the latent diffusion design, our model allows for shape generation with a much higher resolution than previous work (e.g., Matthias [24]). In particular, our method achieves $96^{3}$ resolution which is sufficient for representing rich geometric details of most common-seen objects. Furthermore, we integrate cross- and self-attention mechanisms into the part-aware decoder to establish robust dependencies between the part code and the voxel field. This integration aids in modeling the inherent relationships among various parts, significantly enhancing the precision of part segmentation.
@@
The diffusion model admits multi-modal prompt for shape generation. For example, we have implemented image-conditioned diffusion enabling single-image guided generation. Concretely, a convolutional encoder is employed to extract features from the reference image. Subsequently, these obtained features are combined with the multi-scale features of the denoising UNet. As a result, the image information can effectively guide the shape generation.
@@
To make a fair comparison, we have implemented a supervised version of Part-NeRF. Similar to acquiring color, we use MLP layers to implicitly learn the part emission of 3D position, obtaining the semantic part map via volume rendering. We conduct extensive experiments and compare to the state-of-the-art methods on four classes of common-seen objects. The results show that our method achieves the best FID metric on all classes of data and higher-quality part-aware generation.
2 Related Work
GANs for Shape Generation. Recent developments in generative models have enabled the synthesis of 3D shapes through a variety of representations, including point clouds [1, 35], voxels [25, 14, 26], meshes [10], and neural fields [4, 38, 34, 7, 20]. In particular, the neural field representation has demonstrated significant potential as it enables photo-realistic renderings of 3D objects and can be trained with only 2D image supervision. With the success of Generative Adversarial Network (GAN) [11] in 2D vision, most studies apply it to neural field generation. Pi-GAN [4] and GRAF [33] propose to generate neural fields with stochastic latent codes, and the following GIRAFFE [27] represents the scene as the compositional neural feature field with a similar way. StyleNeRF [12] attempts to increase the generative consistency via the new upsampler and regularization loss. Furthermore, EG3D [3] proposes to use a convolutional network to generate neural tri-plain fields, improving the training efficiency. Different from these methods, we aim to generate part-aware neural voxel field with diffusion models.
@@ 2 Related Work
Diffusion Models for Generating Neural Fields. Recently, the diffusion models [39, 16] has outperformed GANs in various of 2D tasks [8], which inspires several diffusion-based methods to generate neural fields. DreamFusion [29] utilizes the 2D pretrained diffusion model as the distillation model to guide the generation of neural fields for the first time. Following this idea, a series of researches [30, 22, 47] can edit and generate neural fields via text prompt. However, they need to retrain the model for each generation, leading to a time-consuming generating process. Differently, another line of work directly apply the diffusion model to synthesize the neural fields without 2D pretrained diffusion model. DiffRF [24] applies the diffusion model to generate neural voxel fields, however, the 3D convolutional network takes up too much memory cost, limiting the generated resolution. SSD-NeRF [31] proposes a single-stage training paradigm to generate tri-plain fields with 2D convolutional network. In this paper, we propose a 3D latent diffusion model to directly generate neural voxel fields, enabling high-resolution and high-quality generation.
@@
Part-aware Shape Generation. Previous methods [19, 28, 45, 15, 13] usually generate part-aware shapes with point cloud representation, however, they usually use the 3D supervision. Recent PartNeRF [42] proposes to represent the different parts of a shape with different neural radiance fields, however, the divided parts are meaningless and multiple neural fields require much high computational resources. Differently, we carefully design the part relation module and supervise the model with 2D part mask, achieving precise and high-quality part-aware shape generation.
3 Proposed Method
@@
In this section, we introduce a latent 3D diffusion model to generate high-resolution part-aware neural voxel fields. As illustrated in Fig. 1, our approach consists of a latent 3D diffusion model and a part-aware shape decoder. Specifically, the latent 3D diffusion model compresses the object fields into the latent space with much smaller spacial resolution, and then adopts the 3D Unet to conduct the diffusion process. This implementation facilitates the high resolution of the voxel field, allowing the high-quality rendering results. On the other hand, the part-aware shape decoder integrates the part codes into the voxel field, establishing relationships between different parts of the voxel field, and produces RGB images and semantic part maps through ray casting.
3.1 Preliminary
3.1.1 Neural Voxel Fields
@@
Neural radiance fields (NeRFs) [23] represent 3D object implicitly and map the 3D position $\boldsymbol{x}$ and view direction $\boldsymbol{d}$ to the density $\sigma$ and color emission $\boldsymbol{c}$, obtaining the novel view image via volume rendering. Based on NeRFs, DVGO [41] learns neural voxel fields to represent the object density and color. Specifically, there are a density voxel field and a feature voxel field, which are represented by $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$. We can acquire the the density and color feature of position $\boldsymbol{x}$ by querying from $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$ directly, and the color feature and view direction are feed into multilayer perceptron (MLP) layers to get the color emission $\boldsymbol{c}$.
@@
where $M$ is the number of sampled points along the ray, $\alpha_{i}$ is the probability of the termination at point $i$, $\delta_{i}$ is the distance between the sampled point and its adjacent point. In this way, the neural voxel field can be optimized by minimizing the mean square error of the rendered RGB $\hat{\boldsymbol{c}}(\boldsymbol{r})$ and the ground truth color $\boldsymbol{g}_{\boldsymbol{c}}(\boldsymbol{r})$. After the optimization, we can represent the object with $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$. To simply formulation, we concatenate $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$ along the first dimension, resulting in $\mathbf{V}^{\boldsymbol{f}}$ as representation of the neural voxel field.
3.1.2 Decoupled Diffusion Models
Diffusion models has shown the great potential in generative tasks, however, it usually suffers from prolonged inference time. Recent decoupled diffusion models [17] propose to decouple the diffusion process to speed up the sampling process. The decoupled forward diffusion process is controlled by the analytic data attenuation and the standard Wiener process:
@@ 3.1.2 Decoupled Diffusion Models
$q(\mathbf{V}^{\boldsymbol{f}}_{t-\Delta t}|\mathbf{V}^{\boldsymbol{f}}_{t},\mathbf{V}^{\boldsymbol{f}}_{0})=\mathcal{N}(\mathbf{V}^{\boldsymbol{f}}_{t}-\int_{t-\Delta t}^{t}{\boldsymbol{a}_{t}\mathrm{d}t}-\frac{\Delta t}{\sqrt{t}}\boldsymbol{\epsilon},\frac{\Delta t(t-\Delta t)}{t}\mathbf{I}),$ (4)
where $\Delta t$ is the time interval, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the standard normal noise. Benefiting from the analytic attenuation function, we can solve the reversed process with arbitrary time interval via analytic integration, improving the sampling efficiency greatly. In practice, we choose the particular constant form of $\boldsymbol{a}_{t}$: $\boldsymbol{a}_{t}=-\mathbf{V}^{\boldsymbol{f}}_{0}$. In this way, the model is supervised by $\boldsymbol{\epsilon}$ and $\mathbf{V}^{\boldsymbol{f}}_{0}$ simultaneously, and the training objective is formulated by:
@@
3.2 Latent 3D Diffusion Model
@@
3D Unet vs. 2D UNet. Actually, though the latent code $\mathbf{L}_{t}$ is a 4D tensor, it is possible to apply the 2D convolutional operator to this type of data. In practice, we can reshape $\mathbf{L}_{t}$ into a 3D tensor, i.e. $\text{reshape}(\mathbb{R}^{C\times X\times Y\times Z})\rightarrow\mathbb{R}^{CX\times Y\times Z}$, thus adopting 2D UNet to model the denoising process. However, we found that it failed to obtain satisfactory results as shown in Fig. 2. We assume that the 2D UNet architecture hardly learns the 3D consistency of the voxel field, resulting in a perturbed result. In contrast, the 3D UNet architecture provides reasonable performance, demonstrating the necessity of adopting 3D UNet for denoising the neural voxel field.
3.3 Part-aware Shape Decoder
After the denoising process of the latent 3D diffusion model, we can have a neural voxel field containing the density and color feature. However, it can not provide the structural information and rendering results. Therefore, we propose a part-aware shape decoder to generate part-aware details and rendering results.
As shown in Fig. 1, to model the part-aware information, the shape decoder integrates the part code into the voxel field and utilize the part attention to build the relations between different parts of the voxel field. Specifically, the part code is a learnable embedding and can be represented by $\boldsymbol{z}\in\mathbb{R}^{K\times D}$, where $K$ is the number of parts, $D$ is the dimension of the embedding. To incorporate the part code into the neural field, we first project the reconstructed voxel field $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ to have the same dimension as $\boldsymbol{z}$, and then flatten $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ to be a 2D tensor like $\mathbb{R}^{XYZ\times D}$. Thus, we can utilize the multi-head cross-attention [43] to build strong dependency between the part code and the voxel field. Moreover, to divide the voxel field accurately, we employ a self-attention to model the long-range relations between the different parts of the voxel field. After the cross-attention and self-attention, we reshape the tensor back to be a 4D voxel field. We formulate this process by the following equation:

### v3

@@ Abstract
This paper introduces a novel latent 3D diffusion model for generating neural voxel fields with precise part-aware structures and high-quality textures. In comparison to existing methods, this approach incorporates two key designs to guarantee high-quality and accurate part-aware generation. On one hand, we introduce a latent 3D diffusion process for neural voxel fields, allowing generation at significantly higher resolutions to capture rich textural and geometric details accurately. On the other hand, a part-aware shape decoder is introduced to integrate the part codes into the neural voxel fields, guiding accurate part decomposition and producing high-quality rendering results. Importantly, part-aware learning establishes structural relationships to generate texture information for similar regions, thereby facilitating high-quality rendering results. We evaluate our approach across eight different data classes through extensive experimentation and comparisons with state-of-the-art methods. The results demonstrate that our proposed method has superior generative capabilities in part-aware shape generation, outperforming existing state-of-the-art methods. Moreover, we have conducted image- and text-guided shape generation via the conditioned diffusion process, showcasing the advanced potential in multi-modal guided shape generation.
Index Terms:
3D diffusion models, shape generation, part-aware generation.
I Introduction
@@
Recent generative models have emerged for generating 3D shapes with different representations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], among which neural field representation [12, 13, 14, 15] receives increasing attention due to many advantages such as integrated texture and shape generation. Despite the fast and notable progress made along this line of research, what has been largely missed is part/structure awareness. Most existing generative models based on neural fields are part-oblivious. They tend to generate 3D shapes in a holistic manner, without comprehending their compositional parts explicitly.
@@
Generating 3D shapes with part information facilitates several downstream processing/tasks, such as editing, mix-and-match modeling, and segmentation learning. SPAGHETTI [16] and Part-NeRF [17] are two excellent part-aware generative models for 3D shape synthesis with neural field representation. Despite their strengths, there is still potential for enhancing the generative ability. SPAGHETTI relies on 3D supervision and mainly considers the geometry, ignoring the textured shape generation. On the other hand, Part-NeRF requires only 2D supervision but adopts a complex architecture with multiple neural radiance fields, leading to increased complexity in learning and subpar rendering outcomes. Moreover, both models perform object segmentation in an unsupervised manner, resulting in structurally insignificant information.
Figure 1: Our approach involves a latent 3D diffusion process and a part-aware decoding module. The latent 3D diffusion process facilitates the generation of high-resolution voxel fields, while the part-aware decoding module ensures precise part decomposition. Moreover, we admit multi-modal inputs for conditioned shape generation.
@@
We propose a latent 3D diffusion model to generate neural voxel fields with accurate structural details in this paper. As shown in Fig 1, our network is composed of two parts, a latent 3D diffusion model of neural voxel fields and a part-aware shape decoder. Specifically, the latent 3D diffusion model utilizes the advantage of 3D convolution and employs the 3D denoising process in a latent space with a much lower resolution. On the other hand, the part-aware shape decoder incorporates the part code with the neural voxel field to learn part-aware relationships, guiding the precise generation of part-aware shapes and synthesizing high-quality rendering results.
@@
Benefiting from the latent diffusion design, our model allows for shape generation with a much higher resolution than previous work (e.g., Matthias [18] and [19]). In particular, our method achieves $96^{3}$ resolution which is sufficient for representing rich geometric details of most common-seen objects. Furthermore, we integrate cross- and self-attention mechanisms into the part-aware decoder to establish robust dependencies between the part code and the voxel field. This integration aids in modeling the inherent relationships among various parts, significantly enhancing the precision of part segmentation. More importantly, learning part-aware information enhance the correlations between different regions, facilitating the geometry reconstruction and texture modeling in similar parts.
@@
The diffusion model admits multi-modal prompt for shape generation. For example, we have implemented image-conditioned diffusion enabling single-image guided generation. Concretely, a convolutional encoder is employed to extract features from the reference image. Subsequently, these obtained features are combined with the multi-scale features of the denoising UNet. As a result, the image information can effectively guide the shape generation. Moreover, we perform text-conditioned diffusion for text-guided shape generation. We utilize the general language model CLIP [20] to extract embedding for the text input, and perform cross-attention between the text embedding and the multi-scale features of the denoising UNet.
@@
To make a fair comparison, we have implemented a supervised version of Part-NeRF. Similar to acquiring color, we use MLP layers to implicitly learn the part class probability of 3D position, obtaining the semantic part map via volume rendering. We conduct extensive experiments and compare to the state-of-the-art methods on eight classes of common-seen objects. The results show that our method achieves the best metrics (FID, MMD, COV) on all classes of data and higher-quality part-aware generation.
II Related Work
Diffusion Models for Shape Generation. Recent developments in diffusion models [21] have enabled the synthesis of 3D shapes through a variety of representations, including point cloud [22, 23], SDF [24], mesh [1], and neural field [12, 25, 26, 27, 28]. [29] proposes to utilize diffusion models to generate a compact wavelet representation with a pair of coarse and detail coefficient volumes. 3DShape2VecSet [30] introduces a novel representation for neural fields designed for generative diffusion models, which can encode 3D meshes or point clouds as neural fields. DiffFacto [31] learns the distribution of point clouds with part-level control including part-level sampling, mixing and interpolation. EXIM [32] use the text prompt as guidance, generating corresponding truncated signed distance field with diffusion models. However, most of above methods must adopt 3D input data as supervision, which is not available easily. Differently, this paper focus on generating neural voxel fields with differentiable rendering, enabling multi-view image as supervision.
@@ II Related Work
Diffusion Models for Generating Neural Fields. Recently, the diffusion models [33, 21] have outperformed GANs in various 2D tasks [34, 35], which has inspired several diffusion-based methods to generate neural fields. DreamFusion [36] utilizes the 2D pretrained diffusion model as the distillation model to guide the generation of neural fields for the first time. Following this idea, a series of researches [37, 38, 39] can edit and generate neural fields via text prompt. However, they need to retrain the model for each generation, leading to a time-consuming generation process. Differently, another line of work directly applies the diffusion model to synthesize the neural fields without 2D pretrained diffusion models. DiffRF [18] applies the diffusion model to generate neural voxel fields, however, the 3D convolutional network takes up too much memory cost, limiting the generated resolution. SSD-NeRF [40] proposes a single-stage training paradigm to generate tri-plain fields with 2D convolutional network. In this paper, we propose a 3D latent diffusion model to directly generate neural voxel fields, enabling high-resolution and high-quality generation.
@@
Part-aware Shape Generation. Previous methods [41, 42, 43, 16, 44, 45, 46] usually generate part-aware shapes with point cloud representation, however, they typically rely on 3D supervision. Recent PartNeRF [17] proposes to represent the different parts of a shape with different neural radiance fields, however, the divided parts are meaningless and multiple neural fields require much high computational resources. Differently, we carefully design the part relation module and supervise the model with 2D part mask, achieving precise and high-quality part-aware shape generation.
III Proposed Method
@@
In this section, we introduce a latent 3D diffusion model to generate high-resolution part-aware neural voxel fields. As illustrated in Fig. 2, our approach consists of a latent 3D diffusion model and a part-aware shape decoder. Specifically, the latent 3D diffusion model compresses the object fields into the latent space with a much smaller spatial resolution, and then adopts the 3D UNet to conduct the diffusion process. This implementation facilitates the high resolution of the voxel field, allowing the high-quality rendering results. On the other hand, the part-aware shape decoder integrates the part codes into the voxel field, establishing relationships between different parts of the voxel field, and produces RGB images and semantic part maps through ray casting.
III-A Preliminary
III-A1 Neural Voxel Fields
@@
Neural radiance fields (NeRFs) [47] represent 3D object implicitly and map the 3D position $\boldsymbol{x}$ and view direction $\boldsymbol{d}$ to the density $\sigma$ and color emission $\boldsymbol{c}$, obtaining the novel view image via volume rendering. Based on NeRFs, DVGO [48] learns neural voxel fields to represent the object’s density and color. Specifically, there are a density voxel field and a feature voxel field, which are represented by $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$. We can acquire the density and color feature of position $\boldsymbol{x}$ by querying from $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$ directly, and the color feature and view direction are feed into multilayer perceptron (MLP) layers to get the color emission $\boldsymbol{c}$.
@@
where $M$ is the number of sampled points along the ray, $\alpha_{i}$ is the probability of the termination at point $i$, and $\delta_{i}$ is the distance between the sampled point and its adjacent point. In this way, the neural voxel field can be optimized by minimizing the mean square error of the rendered RGB $\hat{\boldsymbol{c}}(\boldsymbol{r})$ and the ground truth color $\boldsymbol{g}_{\boldsymbol{c}}(\boldsymbol{r})$. After the optimization, we can represent the object with $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$. To simplify the formulation, we concatenate $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$ along the first dimension, resulting in $\mathbf{V}^{\boldsymbol{f}}$ as the representation of the neural voxel field.
III-A2 Decoupled Diffusion Models
Diffusion models have shown great potential in generative tasks, however, they usually suffer from prolonged inference time. Recent decoupled diffusion models [49] propose to decouple the diffusion process to speed up the sampling process. The decoupled forward diffusion process is controlled by the analytic data attenuation and the standard Wiener process:
@@ III-A2 Decoupled Diffusion Models
$q(\mathbf{V}^{\boldsymbol{f}}_{t-\Delta t}|\mathbf{V}^{\boldsymbol{f}}_{t},\mathbf{V}^{\boldsymbol{f}}_{0})=\mathcal{N}(\mathbf{V}^{\boldsymbol{f}}_{t}-$ $\int_{t-\Delta t}^{t}{\boldsymbol{a}_{t}\mathrm{d}t}-\frac{\Delta t}{\sqrt{t}}\boldsymbol{\epsilon},$ (4)
$\frac{\Delta t(t-\Delta t)}{t}\mathbf{I}),$
where $\Delta t$ is the time interval, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the standard normal noise. Benefiting from the analytic attenuation function, we can solve the reversed process with arbitrary time intervals via analytic integration, improving the sampling efficiency greatly. In practice, we choose the particular constant form of $\boldsymbol{a}_{t}$: $\boldsymbol{a}_{t}=-\mathbf{V}^{\boldsymbol{f}}_{0}$. In this way, the model is supervised by $\boldsymbol{\epsilon}$ and $\mathbf{V}^{\boldsymbol{f}}_{0}$ simultaneously, and the training objective is formulated by:
@@
III-B Latent 3D Diffusion Model
@@
For conditioned generation, we utilize the classifier-free [52] manner to synthesize shapes corresponding to the conditioned input. Specifically, the conditioned input is fed into a feature extractor, resulting in a feature embedding. We adopt the feature embedding as an additional input of the 3D UNet, thus we obtain the $\mathbf{L}_{\boldsymbol{\phi}}$ and $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}$ by: $\mathbf{L}_{\boldsymbol{\phi}},\boldsymbol{\epsilon}_{\boldsymbol{\phi}}=\text{UNet}_{\boldsymbol{\phi}}(\mathbf{L}_{t},t,\mathbf{e})$, where $\mathbf{e}$ is the feature embedding of the conditioned input.
3D UNet vs. 2D UNet. Actually, though the latent code $\mathbf{L}_{t}$ is a 4D tensor, it is possible to apply the 2D convolutional operator to this type of data. In practice, we can reshape $\mathbf{L}_{t}$ into a 3D tensor, i.e. $\text{reshape}(\mathbb{R}^{C\times X\times Y\times Z})\rightarrow\mathbb{R}^{CX\times Y\times Z}$, thus adopting 2D UNet to model the denoising process. However, we found that the loss curve did not converge and it failed to obtain satisfactory results (details can be found in Sec. IV-C). We assume that the 2D UNet architecture hardly learns the 3D consistency of the voxel field, resulting in a perturbed result. In contrast, the 3D UNet architecture provides reasonable performance, demonstrating the necessity of adopting 3D UNet for denoising the neural voxel field.
TABLE I: Quantitative comparisons to state-of-the-art methods on generative rendering performances. Note the KID metrics are multiplied by 1000.
Method Chair Table Airplane Car Guitar Lamp Laptop Pistol
[…]
Part-NeRF 69.44 14.29 75.78 26.91 83.84 42.10 133.2 77.73 123.16 79.78 106.44 71.09 237.02 203.34 137.4 90.57
EG3D 104.24 36.71 116.88 76.47 108.82 64.77 162.79 128.52 111.83 67.85 131.52 70.24 232.65 198.93 126.33 80.00
DiffRF 43.92 7.25 56.28 22.61 53.33 19.85 88.33 26.65 80.55 49.25 99.77 43.48 198.32 81.59 65.83 45.19
Ours 27.70 5.19 44.27 21.74 36.01 21.10 62.27 29.66 62.79 30.70 75.31 17.43 146.17 107.64 60.58 34.23
III-C Part-aware Shape Decoder
After the denoising process of the latent 3D diffusion model, we can have a neural voxel field containing the density and color features. However, it can not provide the structural information and rendering results. Therefore, we propose a part-aware shape decoder to generate part-aware details and rendering results.
As shown in Fig. 2, to model the part-aware information, the shape decoder integrates the part code into the voxel field and utilizes the part attention to build the relations between different parts of the voxel field. Specifically, the part code is a learnable embedding and can be represented by $\boldsymbol{z}\in\mathbb{R}^{K\times D}$, where $K$ is the number of parts, $D$ is the dimension of the embedding. To incorporate the part code into the neural field, we first project the reconstructed voxel field $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ to have the same dimension as $\boldsymbol{z}$, and then flatten $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ to be a 2D tensor like $\mathbb{R}^{XYZ\times D}$. Thus, we can utilize the multi-head cross-attention [53] to build a strong dependency between the part code and the voxel field. Moreover, to divide the voxel field accurately, we employ a self-attention to model the long-range relations between the different parts of the voxel field. After the cross-attention and self-attention, we reshape the tensor back to be a 4D voxel field. We formulate this process by the following equation: